OpenStudy (youngna):

What is the value of y? 3(y+5)=–2y3(y+5)=–2y A. -5 B. -3 C. 3 D. 5

OpenStudy (dmbroussard1978):

3y+15=-2y -3y -3y 15=-5y -5=-5 -3=y so y=-3

OpenStudy (youngna):

thank you

OpenStudy (dmbroussard1978):

Welcome